OpenStudy (anonymous):

1. 4/5 x - 3 3/4 x = 2 1/2 2. Q - 7 5/6 = 6 1/2 3. Find the product. Simplify if possible -1/3 (8 1/7) just in case you dont understand negative 1 over 3 ( 8 1 over 7) not 81 over 7 little help would be nice

jhonyy9 (jhonyy9):

1. 4x/5 - 15x/4 = 5/2 16x - 75x 50 ---------- = ---- 20 20 - 59x =50 x= -50/59

OpenStudy (callisto):

2. Q - 7 5/6 = 6 1/2 Q - (7x6+5)/6 = (6x2+1)/2 Q - 47/6 = 13/2 Q= 13/2 + 47/6 = 43/3 = 14 1/3

OpenStudy (anonymous):

yea thanks everyone but i could never figure out number 3 just made my head hurt trying

jhonyy9 (jhonyy9):

Q -47/6 =13/2 Q = 13/2 + 47/6 Q = (39+47)/6 Q = 86/6 = 43/3

OpenStudy (callisto):

3. -1/3 (8 1/7) = (-1/3) [(8x7+1)/7] = (-1/3) (57/7) = (-1)(57)/[(3x7)] = -57 / 21 = -2 5/7

OpenStudy (anonymous):

5/7 is not the answer it does not show that

OpenStudy (anonymous):

nice work